Solve equation:
First problem: 4g^2-32g=0
Second problem: 3m^2=-6m

Answer by nerdybill(7384)

First problem:
4g^2-32g=0
factor out common factor of 4g:
(4g)(g-8)=0
setting each term to zero:
g = {0, 8}
.
Second problem:
3m^2=-6m
3m^2+6m = 0
factor out common factor of 3m:
(3m)(m+2) = 0
setting each term to zero:
m = {0, -2}